OpenStudy (anonymous):

Find dy/dx such that x=pi/3 a) csc(pi-x) b) 2^tanx

OpenStudy (gorv):

d(cos(pi-x))/dx = -sin(pi-x) * d(-x)/dx =-sin(pi-x)*(-1) =sin(pi-x) at x= pi/3 sin(pi-pi/3)=(3^1/2)/2

OpenStudy (gorv):

\[y=2^(tanx)\]

OpenStudy (gorv):

take log both side log y=tanx log2 dfferentiate (1/y )*dy/dx=sec^2x * ln2 dy/dx=y *sec^2x * ln2 but y= 2^tanx dy/dx=(2^tanx)*sec^2x * ln2 at x=pi/3 dy/dx = \[2^\sqrt{3}*2^2 * \ln2\]

OpenStudy (gorv):

\[2.772*2^\sqrt{3}\]